# Congruent modular forms and anticyclotomic Iwasawa theory

### v1

@@ Abstract
Let $p$ be an odd prime. Consider normalized newforms $f_{1},f_{2}$ that both satisfy the Heegner hypothesis for an imaginary quadratic field $K$ and suppose that they induce isomorphic residual Galois representations. In the work of Greenberg-Vatsal [1] and Emerton-Pollack-Weston [2], the authors compare the cyclotomic Iwasawa $\mu$ and $\lambda$-invariants of $f_{1}$ and $f_{2}$. We extend this to the anticyclotomic indefinite setting by comparing the BDP $p$-adic $L$-functions attached to $f_{1}$ and $f_{2}$. Using this comparison, we obtain arithmetic implications for both generalized Heegner cycles and the Iwasawa main conjecture.
@@ 1 Introduction
Let $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$ be a normalized newform of weight $2r$ and level $N$ that is an eigenform for all Hecke operators. Fix an odd prime $p\nmid N$. We may attach to $f$ a $p$-adic Galois representation $V_{f}$ and let
@@
be its self-dual Artin twist. We will denote by $\overline{\rho}_{f}$ the associated semisimplified residual representation.
Let $K/\mathbb{Q}$ be an imaginary quadratic field of discriminant $-D_{K}$ and let $p>2$ be a rational prime that split in $K$ as $(p)=\mathfrak{p}\overline{\mathfrak{p}}$. We define the following hypothesis for $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$
$\begin{cases}p\nmid 2(2r-1)!N\phi(N),\\ \text{every prime }\ell\mid N\text{ is split in $K/\mathbb{Q}$}.\\ \end{cases}$
The second condition is known as the strong Heegner hypothesis. In such a setting, we may define the BDP anticyclotomic $p$-adic $L$-function $\mathscr{L}_{\mathfrak{p}}(f)$ attached to $f$ in the sense of [3, 4, 5]. This paper closely follows [5], whose construction of the $p$-adic $L$-function originates from [4]. This $p$-adic $L$-function is defined as an element of the Iwasawa algebra $\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket$ where $\mathcal{W}$ is a certain finite extension of the completed maximal unramified extension $\widehat{\mathbb{Q}_{p}^{nr}}$ and $\Gamma_{K}^{-}$ is the Galois group of the anticyclotomic extension over $K$.
@@
In the indefinite anticyclotomic setting, congruences between the BDP $p$-adic $L$-functions have been studied in [9] for the weight $2$ case. In this setting, Kriz-Li studied the logarithms of Heegner points twisted by unramified characters which are interpolated by the BDP $p$-adic $L$-functions (see [10, Theorem 3.9]). The results in this paper can be seen as generalizations of [9, 10] to forms of higher weights and generalized Heegner cycles. The techniques in this paper differ from [9] and as a result require fewer hypotheses. Moreover, for modular forms that are residually isomorphic with respect to an arbitrary prime power, we are able to show congruences between their $p$-adic $L$-functions with respect to the same prime power (see Theorem 5.10). In a paper by Castella et al. [11], the authors use congruence methods to acquire new instances of the anticyclotomic Iwasawa main conjecture at Eisenstein primes. Their work can be seen as an extension of [1, Theorem (1.3)] to the BDP $p$-adic $L$-function whereas our work (in particular, Theorem 7.5) extends [1, Theorem (1.4)].
@@
Suppose that $f_{1}\in S_{2r_{1}}(\Gamma_{0}(N_{1}))^{\text{new}},f_{2}\in S_{2r_{2}}(\Gamma_{0}(N_{2}))^{\text{new}}$ are normalized Hecke eigenforms whose coefficients lie in some number field $L$. We denote by $L_{p}$ the completion of $L$ with respect to a fixed prime above $p$. Suppose that the induced semi-simplified $\mod{\varpi}^{m}$ Galois representations $\bar{\rho}_{f_{1}},\bar{\rho}_{f_{2}}:\operatorname{Gal}(\overline{L}/L)\rightarrow GL_{2}(\mathcal{O}_{L_{p}}/\varpi^{m}\mathcal{O}_{L_{p}})$ are isomorphic, where $\varpi$ is the uniformizer of $\mathcal{O}_{L_{p}}$. Let $\mathcal{W}$ be the ring of integers of a finite extension of $\widehat{\mathbb{Q}_{p}^{nr}}$ containing $L$.
@@
Suppose that both $f_{1},f_{2}$ satisfy hypothesis (1) for $K/\mathbb{Q}$. We may write $(N_{1})=\mathfrak{N}_{1}\overline{\mathfrak{N}_{1}}$, $(N_{2})=\mathfrak{N}_{2}\overline{\mathfrak{N}_{2}}$ as ideals in $\mathcal{O}_{K}$. For each prime $\ell\mid N_{1}N_{2}$, let $v\mid\mathfrak{N}_{1}\mathfrak{N}_{2}$ be the corresponding prime above $\ell$. Then the following congruence holds:
@@
where for each $\overline{v}$, $\mathscr{P}_{\overline{v}}(f_{1})$ and $\mathscr{P}_{\overline{v}}(f_{2})$ are defined in Definition 5.7. Moreover, one has the following:
@@
For each number field $F$, the embedding $\iota_{p}$ determines a choice of inclusion $F\subset\mathbb{C}_{p}$, or equivalently a prime in $F$ above $p$. We assume that this choice gives rise to the prime $\mathfrak{p}$ in $K$ that is consistent with the splitting $p\mathcal{O}_{K}=\mathfrak{p}\overline{\mathfrak{p}}$ given in the Introduction. We will denote by $F_{p}$ the completion of $F$ with respect to the prime induced by $\iota_{p}$. We will also denote by $\mathbb{A}_{F}$ the adeles of $F$ and $\widehat{F}$ the finite adeles. Moreover, let $F_{\infty}:=\prod_{v\mid\infty}F_{v}$.
@@ 2 Geometric and $p$-adic modular forms
Let $S$ denote a $p$-adic ring (which is also a $\mathbb{Z}_{p}$-algebra) and $R$ denote some $p$-adic $S$-algebra. For an integer $N$, we let $\mu_{N}$ be the group scheme of the $N$-th roots of unity and let $A[N]$ be the group scheme of the $N$-torsion points of an abelian variety $A$.
Consider the isomorphism classes of triples $[(A,\eta_{N},\omega)_{/R}]_{/\simeq}$, where $A/R$ is an elliptic curve and $\eta_{N}:\mu_{N}\rightarrow A[N]$ is the $\Gamma_{1}(N)$-level structure and $\omega\in H^{0}(A/R,\underline{\Omega}^{1}_{A/R})$ is a differential $1$-form. The functor classifying such triples is representable by an affine scheme $\mathcal{M}_{\Gamma_{1}(N)}$ defined over $\mathbb{Z}[1/6N]$ [12, Theorem 3.1].
@@
([12, Section 3.2.3] For each $S$-algebra $R$, consider the set of all triples $[(A,\eta_{N},\omega)_{/R}]\in\mathcal{M}_{\Gamma_{1}(N)}(R)$. A geometric modular form $f$ of weight $k$ and level $\Gamma_{1}(N)$ over $R$ is a rule assigning to such every triple $(A,\eta,\omega)_{/R}$ a value $f(A,\eta,\omega)\in R$ satisfying the following:
@@
    For any $S$-algebra homomorphism $\phi:R\rightarrow R^{\prime}$, we have
    $f((A,\eta,\omega)\otimes_{R}R^{\prime})\simeq\phi(f(A,\eta,\omega))$
@@
    $f((A,\eta_{N}\circ b,\omega)_{/R})=f(A,\eta_{N},\omega)$ for any $b\in(\mathbb{Z}/N\mathbb{Z})^{\times}$ with the canonical action of $(\mathbb{Z}/N\mathbb{Z})$ on $\mu_{N}$ [4, Section 3.1].
We define the $q$-expansion of $f$ as $f(\text{Tate}(q),\eta_{\text{can}},du/u)\in S\llbracket q\rrbracket$, where $\eta_{\text{can}}:\mu_{N}\oplus\mu_{p^{\infty}}\hookrightarrow\mathbf{G}_{m}\rightarrow\mathbf{G}_{m}/q^{\mathbb{Z}}$ is the canonical level structure, and $u$ is the canonical parameter of $\mathbf{G}_{m}=\text{Spec}(\mathbb{Z}[u,u^{-1}]).$
To define $p$-adic modular forms, we first introduce the Igusa scheme $\text{Ig}(N)/\mathbb{Z}_{p}$, which is the moduli space parametrizing isomorphism classes of elliptic curves with $\Gamma_{1}(Np^{\infty})$-structure. More precisely, for each $\mathbb{Z}_{p}$-algebra $R$, $\text{Ig}(N)(R)$ is the isomorphism classes of tuples $[(A,\eta)_{/R}]_{/\simeq}$ where $A/R$ is an elliptic curve and $\eta=\eta_{N}\oplus\eta_{p}:\mu_{N}\oplus\mu_{p^{\infty}}\hookrightarrow A[N]\oplus A[p^{\infty}]$ is an immersion of group schemes ([12, Section 3.2.7]).
@@
([12, Section 3.2.9], [5, Section 2.1]) Denote by $S_{m}:=S/p^{m}S$. We define the space of $p$-adic modular forms of level $\Gamma_{1}(N)$ over $S$, denoted $V_{p}(\Gamma_{1}(N),S),$ as
@@
where $\widehat{\text{Ig}}(N)$ is the formal completion of $\text{Ig}(N)$. In particular, $f$ is a function assigning to each $[(A,\eta)_{/R}]\in\text{Ig}(N)(R)$ a value $f(A,\eta)\in R$, and they satisfy the following conditions:
@@
    For any level structure $\eta_{N}$ of type $\Gamma_{1}(N)$ on the Tate curve $\text{Tate}(q)$, $f(\text{Tate}(q),\eta_{N}\oplus\eta_{p}^{\text{can}})\in S\llbracket q^{1/N}\rrbracket$, where $\eta_{p}^{\text{can}}$ is determined by the canonical image of $\zeta_{p}$ via $\boldmath{\mathbf{G}}_{m}\rightarrow\text{Tate}(q).$
@@
A geometric modular form gives rise to a $p$-adic modular form in the sense of [14, (1.10.15)]: Let $R$ be a complete local $S$-algebra, and let $[(A,\eta)_{/R}]\in\text{Ig}(N)(R)$. The $\Gamma_{1}(Np^{\infty})$-level structure $\eta=\eta_{N}\oplus\eta_{p}$ determines a map $\widehat{\eta}_{p}:\widehat{\mathbf{G}}_{m}\xrightarrow{\sim}\widehat{A}$ [14, (1.10.1)](see also [15, Proposition 1]). This in turn defines a differential $\omega(\widehat{\eta}_{p}):\text{Lie}(A)\simeq\text{Lie}(\widehat{A})\rightarrow\text{Lie}(\widehat{\mathbf{G}}_{m})=R$. One can then define the $p$-adic avatar $\widehat{f}$ of $f$ ([5]) by letting $\widehat{f}(A,\eta)=f(A,\eta,\omega(\widehat{\eta_{p}})).$
@@ 3 CM points
This section follows [5, Section 2]. Let $K$ be an imaginary quadratic field of discriminant $-D_{K}<0$, and suppose that $p$ is split as $p=\mathfrak{p}\overline{\mathfrak{p}}$ in $\mathcal{O}_{K}$. Let $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$ be a newform satisfying hypothesis (1). We may write $N=\mathfrak{N}\overline{\mathfrak{N}}$ for some ideal $\mathfrak{N}$ in $\mathcal{O}_{K}$. For a positive integer $c$, let $\mathcal{O}_{c}:=\mathbb{Z}+c\mathcal{O}_{K}$ be the order of conductor $c$ in $K$, so that $\operatorname{Gal}(K[c]/K)\simeq\mathcal{C}\ell\left(\mathcal{O}_{c}\right)$.
For each prime-to-$\mathfrak{N}\mathfrak{p}$ integral ideal $\mathfrak{a}$ of $\mathcal{O}_{c}$, we may attach a CM point $x_{\mathfrak{a}}=(A_{\mathfrak{a}},\eta_{\mathfrak{a}})$. Such a point is defined over a discrete valuation ring inside $\mathcal{V}=\iota_{p}^{-1}(\mathcal{O}_{\mathbb{C}_{p}})\cap K^{ab}$. If $\mathfrak{a}=\mathcal{O}_{c}$, we write $(A_{c},\eta_{c})$ for $(A_{\mathcal{O}_{c}},\eta_{\mathcal{O}_{c}})$. In this case, we see immediately that $A_{\mathfrak{a}}=A_{c}/A_{c}[\mathfrak{a}]$ and the isogeny $\lambda_{\mathfrak{a}}:A_{c}\rightarrow A_{\mathfrak{a}}$ induced by the quotient map $\mathbb{C}/\mathcal{O}_{c}\rightarrow\mathbb{C}/\mathfrak{a}^{-1}$ yields $\eta_{\mathfrak{a}}=\lambda_{\mathfrak{a}}\circ\eta_{c}.$ An equivalent construction is also available in [4, Section 5.1].
If we let $\mathbb{H}$ be the complex upperhalf plane, then is a complex uniformization
@@
×$\mathbb{H}\times\text{GL}_{2}(\widehat{\mathbb{Q}})$ $\rightarrow\text{Ig}(N)(\mathbb{C})$
@@
where $(A_{x},\eta_{x})$ is the corresponding moduli description. We refer readers to [5, Section 2.1] for the explicit form of this map. Moreover, we will also denote the right action of $GL_{2}(\widehat{\mathbb{Q}})$ on $x=[(\tau_{x},g_{x})]\in\text{Ig}(N)(\mathbb{C})$ as
@@
For the choice of $\vartheta$ given in [5, Section 2.3]:
@@
the embedding $\rho:K\hookrightarrow GL_{2}(\mathbb{Q})$ is of the form
@@
Tensoring with $\mathbb{A}^{(\infty)}_{\mathbb{Q}}$ gives an embedding $\rho:K^{\times}\backslash\widehat{K}^{\times}\hookrightarrow\text{GL}_{2}(\mathbb{Q})\backslash\text{GL}_{2}(\widehat{\mathbb{Q}}).$ Denote by $[\eta,g]$ the image of $(\eta,g)$ under the projection ×$\mathbb{H}\times\text{GL}_{2}(\widehat{\mathbb{Q}})\rightarrow\text{Ig}(N)(\mathbb{C})$. Then $[\eta,g]\in\text{Ig}(N)(K^{ab}),$ and Shimura’s reciprocity law states that
@@
where $\text{rec}_{K}:K^{\times}\backslash\widehat{K}^{\times}\rightarrow\operatorname{Gal}(K^{ab}/K)$ is the geometrically normalized reciprocity law. We apply this to CM points as follows. Let $[(\vartheta,\xi_{c})]\in\text{Ig}(N)(\mathbb{C})$ be the complex uniformization of the CM point $x_{c}:=[(A_{c},\eta_{c})]$ for some $\xi_{c}\in\text{GL}_{2}(\widehat{\mathbb{Q}})$. For an $\mathcal{O}_{c}$-ideal $\mathfrak{a}$ that is prime to $\mathfrak{N}\mathfrak{p}$, let $x_{\mathfrak{a}}=(A_{\mathfrak{a}},\eta_{a})$ and $a\in\widehat{K}^{(cp)\times}$ be an idele such that $\mathfrak{a}=a\widehat{\mathcal{O}}_{c}\cap K$. Both $x_{\mathfrak{a}}$ and $x_{c}$ are defined over $K[c](\mathfrak{p}^{\infty})$ and $x_{\mathfrak{a}}=[(A_{\mathfrak{a}},\eta_{a})]=[\vartheta,\rho(\overline{a})^{-1}\zeta_{c}]=x_{c}^{\sigma_{\mathfrak{a}}}\in\text{Ig}(N)(K[c](\mathfrak{p}^{\infty}))$ where $\sigma_{\mathfrak{a}}=\text{rec}_{K}(a^{-1})_{\mid K[c](\mathfrak{p}^{\infty})}\in\operatorname{Gal}(K[c](\mathfrak{p}^{\infty})/K)$, following Shimura’s reciprocity law.

### v2

@@ Abstract
Let $p$ be an odd prime. Consider normalized newforms $f_{1},f_{2}$ such that both forms satisfy the Heegner hypothesis for an imaginary quadratic field $K$ and suppose that they induce isomorphic residual Galois representations. In the work of Greenberg-Vatsal [1] and Emerton-Pollack-Weston [2], the authors compare the cyclotomic Iwasawa $\mu$ and $\lambda$-invariants of $f_{1}$ and $f_{2}$. We extend this to the anticyclotomic indefinite setting by comparing the BDP $p$-adic $L$-functions attached to $f_{1}$ and $f_{2}$. Using this comparison, we obtain arithmetic implications for both generalized Heegner cycles and the Iwasawa main conjecture.
@@ 1 Introduction
Let $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$ be a normalized newform of weight $2r$ and level $N$ that is an eigenform for all Hecke operators. Fix an odd prime $p\nmid N$ and let $F$ be a finite extension of $\mathbb{Q}_{p}$ containing the Fourier coefficients of $f$. Let $V_{f}$ be the $p$-adic Galois representation attached to $f$ and let
@@
be its self-dual Artin twist. Here, $G_{\mathbb{Q}}$ denotes the absolute Galois group $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Denote by $\overline{\rho}_{f}$ the associated semisimplified residual representation.
Let $K/\mathbb{Q}$ be an imaginary quadratic field of discriminant $D_{K}$ and let $p>2$ be a rational prime that split in $K$ as $(p)=\mathfrak{p}\overline{\mathfrak{p}}$. Define the following hypothesis for $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$
[…]
The second condition is known as the strong Heegner hypothesis. In such a setting, one may construct the Bertolini-Darmon-Prasanna (BDP) anticyclotomic $p$-adic $L$-function $\mathscr{L}_{\mathfrak{p}}(f)$ attached to $f$ in the sense of [4, 5, 3]. This paper closely follows the work of Castella-Hsieh [3], whose construction of the $p$-adic $L$-function originates from the work of Brakočević [5]. This $p$-adic $L$-function is defined as an element of the Iwasawa algebra $\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket$ where $\mathcal{W}$ is a finite extension of the completed maximal unramified extension $\widehat{\mathbb{Q}_{p}^{nr}}$ over $\mathbb{Q}_{p}$ and $\Gamma_{K}^{-}$ is the Galois group of the anticyclotomic extension over $K$.
@@
In the indefinite anticyclotomic setting, congruences between the BDP $p$-adic $L$-functions have been studied in [9] for the weight $2$ case. In this setting, Kriz-Li studied the logarithms of Heegner points twisted by unramified characters which are interpolated by the BDP $p$-adic $L$-functions (see [10, Theorem 3.9]). The results in this paper can be seen as generalizations of [9, 10] to forms of higher weights and generalized Heegner cycles. The techniques in this paper differ from [9] and the results are proved under fewer hypotheses. Moreover, for modular forms that are residually isomorphic with respect to an arbitrary prime power, we are able to show congruences between their $p$-adic $L$-functions with respect to the same prime power (see Theorem 5.12). In a paper by Castella et al. [11], the authors use congruence methods to acquire new instances of the anticyclotomic Iwasawa main conjecture at Eisenstein primes. Their work can be seen as an extension of [1, Theorem (1.3)] to the BDP $p$-adic $L$-function whereas our work (in particular, Theorem 7.5) extends [1, Theorem (1.4)].
@@
Suppose that $f_{1}\in S_{2r_{1}}(\Gamma_{0}(N_{1}))^{\text{new}},f_{2}\in S_{2r_{2}}(\Gamma_{0}(N_{2}))^{\text{new}}$ are normalized Hecke eigenforms whose coefficients lie in $p$-adic field $F$. Suppose that the induced semi-simplified $\mod{\varpi}^{m}$ Galois representations $\bar{\rho}_{f_{1}},\bar{\rho}_{f_{2}}:G_{\mathbb{Q}}\rightarrow\text{GL}_{2}(\mathcal{O}_{F}/\varpi^{m}\mathcal{O}_{F})$ are isomorphic, where $\varpi$ is the uniformizer of $\mathcal{O}_{F}$. Let $\mathcal{W}$ be the ring of integers of a finite extension of $\widehat{\mathbb{Q}_{p}^{nr}}$ containing $F$.
@@
Suppose that both $f_{1},f_{2}$ satisfy hypothesis (1) for $K/\mathbb{Q}$. One may write $(N_{1})=\mathfrak{N}_{1}\overline{\mathfrak{N}_{1}}$, $(N_{2})=\mathfrak{N}_{2}\overline{\mathfrak{N}_{2}}$ as ideals in $\mathcal{O}_{K}$. For each prime $\ell\mid N_{1}N_{2}$, let $v\mid\mathfrak{N}_{1}\mathfrak{N}_{2}$ be the corresponding prime above $\ell$. Then the following congruence holds:
@@
where $\mathscr{P}_{\overline{v}}(f_{1})$ and $\mathscr{P}_{\overline{v}}(f_{2})$ are defined in Definition 5.9. Moreover, one has the following:
@@
For each number field $L$, the embedding $\iota_{p}$ determines a choice of inclusion $L\subset\mathbb{C}_{p}$, or equivalently a prime in $L$ above $p$. We assume that this choice gives rise to the prime $\mathfrak{p}$ in $K$ that is consistent with the splitting $p\mathcal{O}_{K}=\mathfrak{p}\overline{\mathfrak{p}}$ given in the Introduction. We will denote by $L_{p}$ the completion of $L$ with respect to the prime induced by $\iota_{p}$. We will also denote by $\mathbb{A}_{L}$ the adeles of $L$ and $\widehat{L}$ the finite adeles. Moreover, let $L_{\infty}:=\prod_{v\mid\infty}L_{v}$.
@@ 2 Geometric and $p$-adic modular forms
Let $S$ denote a $\mathbb{Z}_{(p)}$-algebra and let $R$ denote some algebra over $S$. For an integer $N$, let $\mu_{N}$ be the group scheme of the $N$-th roots of unity and let $A[N]$ be the group scheme of the $N$-torsion points of an abelian variety $A$.
Consider the isomorphism classes of triples $(A,\eta_{N},\omega)_{/R}]_{/\simeq}$, where $A/R$ is an elliptic curve and $\eta_{N}:\mu_{N}\rightarrow A[N]$ is the $\Gamma_{1}(N)$-level structure and $\omega\in H^{0}(A/R,\underline{\Omega}^{1}_{A/R})$ is a differential $1$-form. The functor classifying such triples is representable by an affine scheme $\mathcal{M}_{\Gamma_{1}(N)}$ defined over $\mathbb{Z}[1/6N]$ [12, Theorem 3.1].
@@
([12, Section 3.2.3] For each $S$-algebra $R$, consider the set of all triples $(A,\eta_{N},\omega)_{/R}]\in\mathcal{M}_{\Gamma_{1}(N)}(R)$. A geometric modular form $f$ of weight $k$ and level $\Gamma_{1}(N)$ over $R$ is a rule assigning to such every triple $(A,\eta,\omega)_{/R}$ a value $f(A,\eta,\omega)\in R$ satisfying the following:
@@
    For any $S$-algebra homomorphism $\phi:R\rightarrow R^{\prime}$, we have f((A, η, ω) _R R’) ϕ(f(A, η, ω))
@@
    $f((A,\eta_{N}\circ b,\omega)_{/R})=f(A,\eta_{N},\omega)$ for any $b\in(\mathbb{Z}/N\mathbb{Z})^{\times}$ with the canonical action of $(\mathbb{Z}/N\mathbb{Z})$ on $\mu_{N}$ [5, Section 3.1].
Define the $q$-expansion of $f$ as $f(\text{Tate}(q),\eta_{\text{can}},du/u)\in S\llbracket q\rrbracket$, where $\eta_{\text{can}}:\mu_{N}\oplus\mu_{p^{\infty}}\hookrightarrow\mathbf{G}_{m}\rightarrow\mathbf{G}_{m}/q^{\mathbb{Z}}$ is the canonical level structure, and $u$ is the canonical parameter of $\mathbf{G}_{m}=\text{Spec}(\mathbb{Z}[u,u^{-1}]).$
To define $p$-adic modular forms, we first recall the Igusa scheme $\text{Ig}(N)/\mathbb{Z}_{p}$, which is the moduli space parametrizing isomorphism classes of elliptic curves with $\Gamma_{1}(Np^{\infty})$-structure. More precisely, for each $\mathbb{Z}_{p}$-algebra $R$, $\text{Ig}(N)(R)$ is the isomorphism classes of tuples $(A,\eta)_{/R}$ where $A/R$ is an elliptic curve and $\eta=\eta_{N}\oplus\eta_{p}:\mu_{N}\oplus\mu_{p^{\infty}}\hookrightarrow A[N]\oplus A[p^{\infty}]$ is an immersion of group schemes ([12, Section 3.2.7]).
@@
([12, Section 3.2.9], [3, Section 2.1]) Let $S$ be a $p$-adic ring and denote $S_{m}:=S/p^{m}S$. Define the space of $p$-adic modular forms of level $\Gamma_{1}(N)$ over $S$, denoted $V_{p}(\Gamma_{1}(N),S),$ as
@@
where $\widehat{\text{Ig}}(N)$ is the formal completion of $\text{Ig}(N)$. In particular, $f$ is a function assigning to each $(A,\eta)_{/R}]\in\text{Ig}(N)(R)$ a value $f(A,\eta)\in R$, and they satisfy the following conditions:
@@
    For any level structure $\eta_{N}$ of type $\Gamma_{1}(N)$ on the Tate curve $\text{Tate}(q)$, $f(\text{Tate}(q),\eta_{N}\oplus\eta_{p}^{\text{can}})\in S\llbracket q^{1/N}\rrbracket$, where $\eta_{p}^{\text{can}}$ is determined by the canonical image of $\zeta_{p}$ via ${\mathbf{G}}_{m}\rightarrow\text{Tate}(q).$
@@
A geometric modular form gives rise to a $p$-adic modular form in the sense of [14, (1.10.15)]: Let $R$ be a complete local $S$-algebra, and let $(A,\eta)_{/R}]\in\text{Ig}(N)(R)$. The $\Gamma_{1}(Np^{\infty})$-level structure $\eta=\eta_{N}\oplus\eta_{p}$ determines a map $\widehat{\eta}_{p}:\widehat{\mathbf{G}}_{m}\xrightarrow{\sim}\widehat{A}$ [14, (1.10.1)](see also [15, Proposition 1]). This in turn defines a differential $\omega(\widehat{\eta}_{p}):\text{Lie}(A)\simeq\text{Lie}(\widehat{A})\rightarrow\text{Lie}(\widehat{\mathbf{G}}_{m})=R$. One can then define the $p$-adic avatar $\widehat{f}$ of $f$ ([3]) by letting $).\widehat{f}(A,\eta)=f(A,\eta,\omega(\widehat{\eta_{p}}$)).
@@ 3 CM points
This section follows [3, Section 2]. Let $K$ be an imaginary quadratic field of discriminant $D_{K}<0$, and suppose that $p$ is split as $p=\mathfrak{p}\overline{\mathfrak{p}}$ in $\mathcal{O}_{K}$. Let $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$ be a newform satisfying hypothesis (1). One may write $N=\mathfrak{N}\overline{\mathfrak{N}}$ for some ideal $\mathfrak{N}$ in $\mathcal{O}_{K}$. For a positive integer $c$, let $\mathcal{O}_{c}:=\mathbb{Z}+c\mathcal{O}_{K}$ be the order of conductor $c$ in $K$, so that $\operatorname{Gal}(K[c]/K)\simeq\mathcal{C}\ell\left(\mathcal{O}_{c}\right)$.
For each prime-to-$\mathfrak{N}\mathfrak{p}$ integral ideal $\mathfrak{a}$ of $\mathcal{O}_{c}$, there is a CM point $x_{\mathfrak{a}}=(A_{\mathfrak{a}},\eta_{\mathfrak{a}})$ as constructed in [3, Section 2.3] where $A_{\mathfrak{a}}$ is the complex elliptic curve $\mathbb{C}/\mathfrak{a}^{-1}$. Such a point is defined over a discrete valuation ring inside $\mathcal{V}=\iota_{p}^{-1}(\mathcal{O}_{\mathbb{C}_{p}})\cap K^{ab}$. If $\mathfrak{a}=\mathcal{O}_{c}$, we write $(A_{c},\eta_{c})$ for $(A_{\mathcal{O}_{c}},\eta_{\mathcal{O}_{c}})$. In this case, it is immediate that $A_{\mathfrak{a}}=A_{c}/A_{c}[\mathfrak{a}]$ and the isogeny $\lambda_{\mathfrak{a}}:A_{c}\rightarrow A_{\mathfrak{a}}$ induced by the quotient map $\mathbb{C}/\mathcal{O}_{c}\rightarrow\mathbb{C}/\mathfrak{a}^{-1}$ yields $\eta_{\mathfrak{a}}=\lambda_{\mathfrak{a}}\circ\eta_{c}.$ An equivalent construction is also available in [5, Section 5.1].
Let $\widehat{\mathbb{Q}_{p}^{nr}}$ be the $p$-adic completion of maximal unramified extension $\mathbb{Q}_{p}^{nr}$ of $\mathbb{Q}_{p}$. If $\mathfrak{a}$ is a prime-to-$\mathfrak{N}\mathfrak{p}$ ideal of $\mathcal{O}_{c}$ with $p\nmid c$, then $(A_{\mathfrak{a}},\eta_{\mathfrak{a}})$ has a model defined over $\mathcal{V}^{nr}:=\mathcal{W}\cap K^{ab}$ where $\mathcal{W}$ is the ring of integers of $\widehat{\mathbb{Q}_{p}^{nr}}$. We will also denote this model by $(A_{\mathfrak{a}},\eta_{\mathfrak{a}})$ for the rest of this article. For $\mathcal{O}_{c}=\mathcal{O}_{K}$, wr ite $A$ for $A_{\mathcal{O}_{K}}$ and fix a Néron differential $\omega_{A}$ of $A$ over $\mathcal{V}^{nr}$.
If we let $\mathbb{H}$ be the complex upperhalf plane, then there is a complex uniformization
@@
[…]
@@
[…]
@@
For the choice of $\vartheta$ given in [3, Section 2.3]:
@@
the embedding $\rho:K\hookrightarrow\text{GL}_{2}(\mathbb{Q})$ is of the form
@@
[…]
@@
[…]
$x_{\mathfrak{a}}=[(A_{\mathfrak{a}},\eta_{a})]=[\vartheta,\rho(\overline{a})^{-1}\zeta_{c}]=x_{c}^{\sigma_{\mathfrak{a}}}\in\text{Ig}(N)(K[c](\mathfrak{p}^{\infty}))$
where $\sigma_{\mathfrak{a}}=\text{rec}_{K}(a^{-1})_{\mid K[c](\mathfrak{p}^{\infty})}\in\operatorname{Gal}(K[c](\mathfrak{p}^{\infty})/K)$, following Shimura’s reciprocity law.